OpenStudy (anonymous):

Solve for the variable -m/4.2=5

OpenStudy (anonymous):

-m/4.2=5 -m=5(4.2) (multiply 4.2 on both sides) -m=21 (multiply 5 and 4.2) m= -21

OpenStudy (anonymous):

thank you!